# KinesCeTI: A Modular and Size-Adaptable Force Feedback Glove with Interchangeable Actuation for the Index and Thumb

### v1

@@ IV User studies
  • A pick and place manipulation task in VR.
@@
Each study was done with written informed consent of each participant. These studies followed ethical guidelines and were approved by the TU Dresden Ethics Committee, with approval number SR-EK-288072024.
@@ IV-A1 Experiment definition
This experiment evaluated the exoskeleton’s adaptability to different hand sizes and mobility range, along with the impact of its structure (without feedback) on real-object manipulation. Two hypotheses were tested: (1) the exoskeleton adapts to a wide range of hand sizes while preserving natural finger motion and (2) the strap thimble allows effective manipulation of real objects.
The evaluation consisted of four parts. First, the measurement of 25 right-hand parameters (Fig. 9, Table III). A subset of these parameters was compared to anthropometric data from the ANSUR database of 1000 males and 1300 females [mclain2010use]. Second, participants performed 16 hand poses with both the free hand and with the exoskeleton using fingertip and strap thimbles (Fig. 10). Third, a customized Six-Hole Peg test, adapted from the Nine-Hole Peg Test [feys2017nine], was conducted to assess fine-motor manipulation (Fig. 11). Participants performed three repetitions of the task with the free hand and with the exoskeleton attached at the second phalanges using strap thimbles, after a short training period. Number of dropped pegs and executing time were measured. Finally, ergonomic evaluation was performed using NASA-TLX questionnaires [hart2006nasa] for both conditions of the peg test and a customized set of 16 ergonomic questions (Table IV).
TABLE III: Hand Measurement Definitions
ID Description ID Description
1 IN-DP length 15 IN-DP circumference
2 IN-MP length 16 IN-DIP circumference
3 IN-PP length 17 IN-MP circumference
4 IN-MC length 18 IN-PIP circumference
5 IN phalanx-crease length 19 IN-PP circumference
6 MI-MC length 20 TH-DP circumference
7 RI-MC length 21 TH-IP circumference
8 LI-MC length 22 TH-PP circumference
9 TH-DP length 23 Knuckles section circumference
10 TH-PP length 24 Palm section circumference
11 TH-MC length 25 Wrist section circumference
12 Knuckles section length 26 IN length to knuckle*
13 Palm section length 27 IN length to crease*
14 Wrist section length 28 TH length to wrist*
Abbreviations: IN=index, TH=thumb, MI=middle, RI=ring, LI=little, DP=distal phalanx, MP=middle phalanx, PP=proximal phalanx, MC=metacarpal, MCP/PIP/DIP/IP=standard joints. *Derived segments, calculated as in Fig. 9.
Figure 9: Dimensions of the hand, according to Table III.
Figure 10: Hand postures performed in the RoM test.
Figure 11: Setup for the Six-Hole Peg task.
TABLE IV: Customized Ergonomics Questionnaire for the first experiment.
ID Question
1 How well does the glove fit your hand?
2 How well does the tip thimble fit on your index finger?
3 How well does the tip thimble fit on your thumb?
@@
4 How well does the strap thimble fit on the second digit of your index?
@@
5 How well does the strap thimble fit on the second digit of your thumb?
6 How would you rate the weight of the exoskeleton?
7 To what extent did the exoskeleton restrict your motion in the index?
8 To what extent did the exoskeleton restrict your motion in the thumb?
9 To what extent did the exoskeleton restrict overall hand motion?
@@
10 How stable did the exoskeleton feel while moving your fingertip thimbles?
@@
11 How stable did the exoskeleton feel while moving your strap thimbles?
12 How comfortable was the exoskeleton material on your skin?
13 Did your hand feel hot or sweaty while wearing the exoskeleton?
14 To what extent did the exoskeleton affect performance in the Peg Test?
@@
15 How would you compare your speed in the Peg Test with and without the exoskeleton?
16 How would you rate the overall comfort of the exoskeleton?
@@
Scales: 0–100 with anchors specific to each question. Intermediate anchors (25, 75) were also shown (e.g., “Somewhat uncomfortable,” “Mostly stable”) Q1–5: 0 = No fit, 50 = Acceptable, 100 = Perfect; Q6: 0 = Very light, 50 = Neutral, 100 = Very heavy; Q7–9: 0 = No restriction, 50 = Some restriction, 100 = Extreme restriction; Q10–11: 0 = Very unstable, 50 = Neutral, 100 = Fully stable; Q12 & Q16: 0 = Very uncomfortable, 50 = Neutral, 100 = Very comfortable; Q13-14: 0 = No/None, 50 = Moderately, 100 = Extremely; Q15: 0 = Much faster free hand, 50 = Same, 100 = Much faster exo.
@@ IV-A2 Experiment results
The study included 20 participants (19 males, 1 female), aged 24 to 44 years (mean= 32.05). Seventeen participants were right-handed and three left-handed. Table V summarizes the hand measurements obtained in our study. To assess generalizability, these results were compared with anthropometric statistics from the ANSUR database [mclain2010use]. Because some of our measurements were not directly compatible due to different measurement definitions, only the overlapping variables were extracted and compared (Table VI). As our dataset includes mixed genders while ANSUR reports separate male and female distributions, we combined the latter into a unified distribution under the assumption of normality. For each parameter, the percentiles corresponding to our minimum and maximum values were computed, yielding an estimate of the coverage (%) of the ANSUR population. Despite the limited set of comparable parameters, the results indicate that our hand measurements span a substantial portion of the ANSUR size range, with an average coverage of $76.01\pm 10.55\%$ across the compared variables.
TABLE V: Measured hand parameters
ID Mean SD Min Max ID Mean SD Min Max
1 25.04 1.81 21.3 28.58 15 50.5 3.51 44 57
2 25.24 2.8 20.98 32.53 16 54.8 2.8 49 60
3 43.66 3.84 36.47 51.61 17 59.05 2.8 54 64
4 73.38 3.69 66.37 79.66 18 66.5 2.64 62 71
5 25.61 2.35 20.29 30.15 19 69.5 3.15 62 74
6 67.11 4.66 59.1 75.38 20 61.45 3.85 55 69
7 63.07 3.35 57.59 69.13 21 67.25 3.52 61 73
8 63.05 5.43 54.73 76.62 22 69.6 3.29 62 75
9 31.36 2.44 26.6 35.45 23 205.1 11.09 174 221
10 34.91 4.29 28.08 42.3 24 218.65 12.3 189 233
11 50.23 4.5 42.81 59.54 25 174.25 8.73 159 188
12 83.89 3.94 74.32 91.15 26 93.93 7.32 80.31 107.61
13 86.52 4.06 76.97 91.87 27 75.88 5.69 64.34 82.79
14 60.25 3.77 52.01 69.72 28 116.5 8.06 101.88 130.07
Units: mm.
TABLE VI: Common Hand Measurements: Comparison with McLain [mclain2010use]
Combined data from McLain [mclain2010use](mm) Comparison results (%)
ID Mean SD Min Max pMin pMax %
1 26.76 2.62 19 36 1.85 75.64 73.78
2 21.75 2.41 14 32 37.42 100 62.58
9 32.41 3.09 22 45 3.01 83.74 80.72
12 84.24 6.71 66 106 6.97 84.85 77.88
14 60.83 5.86 46 82 6.63 93.54 86.91
16 53.67 3.72 45 65 10.49 95.55 85.06
18 64.33 4.05 56 74 28.26 95.01 66.76
21 67.04 3.53 56 81 12.86 79.73 73.94
23 198.41 16.37 158 247 6.8 91.64 84.81
25 161.36 13.6 130 204 43.12 97.49 54.37
27 72.08 5.51 56 92 8.02 97.4 89.39
% == pMax - pMin.
Figure 12: Motion restrictions for the exoskeleton with each thimble attachment in comparison with the free hand.
@@
In the motion tests with fingertip and strap thimbles, participants were able to perform 12 and 10 out of 16 hand postures, respectively, without restriction (Fig. 12). For the remaining postures, restrictions were observed, differing between the two thimble types. Most limitations occurred near the end of the motion range. With the exception of one participant in each setup who reported moderate restriction for the 11th posture, all others reported only minor restriction. Overall, results indicate that the exoskeleton allows most natural hand postures, though improvements remain possible.
@@
In the Six-Hole Peg test, no failures occurred with the free hand and only four failures in total with the exoskeleton (each by a different participant in the third round). As these were rare events, no statistical analysis was conducted on failure rates. In contrast, execution times differed significantly between conditions. For each subject and condition, the average of the three repetitions was calculated. Since data were not normally distributed, results were analyzed using medians and IQRs across participants. Median task time was $19.99$ s (IQR =15.69=15.6922.7822.78) with the exoskeleton, compared to $14.60$ s (IQR =12.79=12.7916.2116.21) with the free hand. The median paired difference was $4.70$ s, 95% CI [$3.52$, $6.80$]. A one-sided Wilcoxon signed-rank test confirmed longer times with the exoskeleton, $W=210$, $z=3.91$, $p<.001$, with a large effect size ($r=.88$). For descriptive context, mean times were $14.89\pm 2.64$ s for the free hand and $20.48\pm 4.94$ s for the exoskeleton, corresponding to a mean increase of $5.58\pm 2.71$ s, or $36.5\pm 13.5$%. These findings show that use of the exoskeleton without feedback increased execution times and reduced performance efficiency. Nonetheless, aside from the four isolated failures, participants were still able to complete the manipulation task with the exoskeleton.
@@
NASA-TLX scores were compared between free hand and exoskeleton conditions using one-sided Wilcoxon signed-rank tests, under the alternative hypothesis that workload would be higher with the exoskeleton (Fig. 13 and Table VII). After Bonferroni-Holm correction, the exoskeleton condition was rated significantly higher in mental demand, physical demand, effort, frustration and overall workload (all $p<.05$, one-sided). Surprisingly, temporal demand and performance did not show statistically significant differences (p = .89 and .5, respectively), even though objectively measured metrics indicate otherwise.
@@
TABLE VII: NASA-TLX scores of the Six-Hole Peg test reported as median [IQR] and Wilcoxon signed-rank results (n=20).
@@
TABLE VIII: Customized Ergonomics Questionnaire Results (Median [IQR])
ID Mdn [IQR] ID Mdn [IQR]
1 80 [73–91.5] 9 23.5 [9–25.25]
2 78 [67.5–90] 10 75 [48.75–89.25]
3 76.5 [63–89.25] 11 82.5 [75–90.25]
4 85 [76.75–96] 12 67 [50–81]
5 80 [65.75–93.25] 13 2 [0–28.25]
6 61 [48.25–67] 14 38 [14–64.25]
7 10.5 [5.25–25.25] 15 26 [20–35.25]
8 23 [13.75–41.75] 16 70 [50–84.5]
Values are reported as median [IQR] on 0–100 scales, with anchors defined in Table IV.
The ergonomics questionnaire results (Table VIII) can be grouped into several categories. Fit (Q1–5) was rated positively, with medians between 76.5 and 85, indicating generally good adaptation of the glove and thimbles. Weight (Q6) was judged moderately heavy (Mdn = 61, IQR = 48.25–67), with most values between neutral and heavy. Restriction (Q7–9) was low overall: the index finger showed the least limitation (Mdn = 10.5), followed by the thumb (Mdn = 23), while overall hand restriction remained below 50 for almost all participants. Stability (Q10–11) was consistently high, with both fingertip and strap thimbles rated mostly stable (Mdn $\geq$ 75). Comfort (Q12, Q16) was also positive (Mdn = 67–70), although not maximal. Heat and sweating (Q13) were reported as minimal (Mdn = 2, IQR = 0–28.25). Finally, participants acknowledged an impact on perceived performance in the Six-Hole Peg test, rating completion as faster with the free hand (Q14–15, Mdn = 26–38), consistent with the observed execution times.
Overall, this first test evaluated ergonomics, mobility and the effects of wearing the exoskeleton. The device enabled most movements with minimal restriction in a group of subjects that moderately resembled the general population, while offering satisfactory yet improvable ergonomics. With regard to movement and task performance, the Six-Hole Peg task showed increased execution time when using the exoskeleton, but very low impact on effectiveness (low failure rate).
IV-B Pick and place manipulation task
@@ IV-B1 Experiment definition
The second experiment assessed the influence of force feedback from the exoskeleton’s ratchet–pawl mechanism compared to free-hand control in a virtual pick-and-place task. Participants were required to grasp a virtual ball and place it into one of two target rings (Fig. 14(a)). The task was implemented in CHAI3D. The independent variables were:
@@
  • Tolerance threshold for failure with two values (0.025 and 0.03). During object pickup, fingertip distance was recorded and the threshold defined the allowable variation. Smaller thresholds increased the likelihood of dropping or over-squeezing the object (thus higher difficulty). These values were set after pilot testing and are not meant to represent real-world dimensions.
  • Target ring: far (Goal 1) versus near (Goal 2) from the subject, both aligned along the Y-axis (forward from the user) and equidistant from the ball’s initial position on the right.
@@
The experimental hypothesis was that force feedback from the exoskeleton would reduce execution time and failure rates and yield more favorable subjective ratings, compared to free-hand interaction in virtual manipulation. Data were collected at two levels. Objective metrics included execution time per trial and the number of failures. Subjective metrics included a NASA-TLX questionnaire for each hand condition, followed by a questionnaire on force and acoustics perception and overall preference (Table IX).
Figure 14: Virtual setups for (a) the pick an place test and (b) softness discrimination experiment.
TABLE IX: Questionnaire for the Virtual Pick-and-Place Task

### v2

@@ IV User studies
  • A pick-and-place manipulation task in VR.
@@
All studies were conducted with written informed consent and approved by the TU Dresden Ethics Committee (SR-EK-288072024). As the device is designed for right-hand use only, all participants performed the tasks with the right hand, regardless of handedness. For the two VR experiments, a brief automatic software calibration was performed at the beginning, during which participants executed three pinch motions to define the motion range for the virtual digits.
@@ IV-A1 Experiment definition
This experiment evaluated the exoskeleton’s adaptability to different hand sizes and range of motion, as well as the effect of its structure (without feedback) on real-object manipulation. Two hypotheses were tested: (1) the exoskeleton adapts to a wide range of hand sizes while preserving natural digit motion, and (2) the strap thimble enables effective manipulation. For (1), objective metrics included 25 right-hand measurements (Fig. 9, Table III), with a subset compared to ANSUR data [17] to estimate population coverage. Digit motion was evaluated across 16 predefined hand poses under three conditions (free hand, exoskeleton with fingertip and strap thimbles), and assessed qualitatively through restriction levels assigned by the experimenter based on visual estimation of posture completion (Fig. 10(10)). For (2), objective metrics included a customized Six-Hole Peg test adapted from the Nine-Hole Peg Test [28]. Participants performed three repetitions with the free hand and with the exoskeleton attached at the second phalanx, after a short training period. Performance was quantified by completion time, defined as the time required to place all pegs and return them to their original positions, and peg-drop count, defined as pegs falling outside the target holes. Subjective evaluation included NASA-TLX [21] for each condition in the peg test, and a customized set of 16 ergonomic questions (Table IV).
Figure 10: Range-of-motion (RoM) test: (a) Hand postures, and (b) Motion restrictions with each thimble attachment in comparison with the free hand.
Figure 11: Setup for the Six-Hole Peg task: (a) Peg assembly, and (b) Exoskeleton with strap thimbles at the second phalanx.
TABLE IV: Ergonomics Questionnaire and Results (Study 1)
ID Question Mdn [IQR]
1 How well does the glove fit your hand? 80 [73–91.5]
2 How well does the tip thimble fit on your index finger? 78 [67.5–90]
3 How well does the tip thimble fit on your thumb? 76.5 [63–89.25]
@@
4 How well does the strap thimble fit on the second digit of your index? 85 [76.75–96]
@@
5 How well does the strap thimble fit on the second digit of your thumb? 80 [65.75–93.25]
6 How would you rate the weight of the exoskeleton? 61 [48.25–67]
7 To what extent did the exoskeleton restrict your motion in the index? 10.5 [5.25–25.25]
8 To what extent did the exoskeleton restrict your motion in the thumb? 23 [13.75–41.75]
9 To what extent did the exoskeleton restrict overall hand motion? 23.5 [9–25.25]
@@
10 How stable did the exoskeleton feel while moving your fingertip thimbles? 75 [48.75–89.25]
@@
11 How stable did the exoskeleton feel while moving your strap thimbles? 82.5 [75–90.25]
12 How comfortable was the exoskeleton material on your skin? 67 [50–81]
13 Did your hand feel hot or sweaty while wearing the exoskeleton? 2 [0–28.25]
14 To what extent did the exoskeleton affect performance in the Peg Test? 38 [14–64.25]
@@ 12 How comfortable was the exoskeleton material on your skin? 67 [50–81]
15 How would you compare your speed in the Peg Test with and without the exoskeleton? 26 [20–35.25]
16 How would you rate the overall comfort of the exoskeleton? 70 [50–84.5]
@@
Scales: 0–100. Intermediate anchors (25, 75) were also provided. Q1–5 fit (0=no fit, 50=acceptable, 100=perfect); Q6 weight (0=very light, 50=neutral, 100=very heavy); Q7–9 restriction (0=none, 50=some, 100=extreme); Q10–11 stability (0=very unstable, 50=neutral, 100=fully stable); Q12 & Q16 comfort (0=very uncomfortable, 50=neutral, 100=very comfortable); Q13–14 extent (0=none, 50=moderately, 100=extremely); Q15 speed (0=much faster free hand, 50=same, 100=much faster exo).
@@ IV-A2 Experiment results
The study included 20 participants (19 males, 1 female), aged 24 to 44 years (mean = 32.05). Seventeen participants were right-handed and three left-handed. Table III summarizes the hand measurements obtained in our study. To assess generalizability, these results were compared with anthropometric statistics from the ANSUR database [17]. Because some of our measurements were not directly compatible due to different measurement definitions, only the overlapping variables were extracted and compared (Table V). As the sample is predominantly male (95 %), only the male data of both datasets were compared, while the inclusion of a single female participant is acknowledged as a limitation of the sample. As the ANSUR dataset reports percentiles for each parameter (p1–p99), inverse interpolation was used to estimate the percentiles corresponding to the observed minimum and maximum measurements, yielding population coverage (%). Across the compared parameters, the minimum estimated coverage was 52.12%, representing a lower bound and indicating that, for the most restrictive parameter, approximately half of the male ANSUR population is represented by our dataset.
TABLE V: Comparison with Male ANSUR Data [17]
Measured (male, mm) Greiner[17] (male, mm) Comparison results (%)
ID Min Max Mean SD Min Max pMin pMax %
1 22.22 28.58 28.4 2.3 21 36 <1<1 53.12 52.12*
2 20.98 32.53 22.6 2.4 16 32 24.98 >99>99 74.02*
9 26.65 35.45 34.5 2.6 27 45 <1<1 64.26 63.26*
12 79.43 91.15 90.4 4.2 79 106 <1<1 57.09 56.09*
14 55.98 69.72 65.8 4.5 53 82 1.45 80.82 79.36
16 51 60 57.4 1.6 53 65 <1<1 94.79 93.79*
18 62 71 68.4 1.8 64 74 <1<1 92.57 91.57*
21 62 73 72.3 2.9 63 81 <1<1 59.54 58.54*
23 193 221 213.9 9.8 182 247 1.65 76.56 74.91
25 159 188 174.3 8.2 143 204 3.1 95.26 92.16
27 64.34 82.79 75.3 4.9 58 92 1.27 93.68 92.42
% == pMax - pMin. * Lower-bound coverage computed using p1 and p99 when values fall outside the ANSUR percentile range.
@@
In the motion tests, restriction levels across the 16 postures varied with thimble type (Fig. 10(10)). Participants were able to perform 12 and 10 out of 16 hand postures, respectively, without restriction. Limitations differed between thimble types and occurred predominantly near the end of the motion range. With the exception of one participant who reported moderate restriction in posture 11, all restrictions were minor. Slight restrictions were observed in posture 6 due to contact between the second linkage and the thimbles during full thumb radial abduction, and in a small number of cases in posture 8 due to interaction between the strap thimble and the palmar attachment strap. For postures 9–11, a slightly improved range of motion was observed with the strap thimble compared to the fingertip thimble, attributed to the latter being more likely to reach the limits of the exoskeleton motion. Overall, the exoskeleton allows most natural hand postures, with remaining limitations primarily associated with extreme configurations outside typical manipulation gestures (13–16), which were performed without restriction.
@@
In the Six-Hole Peg test, no drops occurred with the free hand, and only four drops were observed with the exoskeleton (out of 60 rounds, 720 pick-and-place actions), each by a different participant in the third round. As drops were rare, no statistical analysis was conducted on drop counts. Execution times, however, differed significantly between conditions. For each subject and condition, times were averaged across three repetitions, and medians and IQRs were used due to non-normality. Median task time was $19.99$ s (IQR =15.69=15.6922.7822.78) with the exoskeleton and $14.60$ s (IQR =12.79=12.7916.2116.21) with the free hand, yielding a median paired difference of $4.70$ s, 95% CI [$3.52$, $6.80$]. A one-sided Wilcoxon signed-rank test confirmed longer times with the exoskeleton, $W=210$, $z=3.91$, $p<.001$, with a large effect size ($r=.88$). For context, mean times were $14.89\pm 2.64$ s (free hand) and $20.48\pm 4.94$ s (exoskeleton), corresponding to a mean increase of $5.58\pm 2.71$ s, or $36.5\pm 13.5$%. During this test, the exoskeleton was used passively, without applying any force feedback. Under this condition, longer execution times were observed compared to the free hand condition. The increased execution times may reflect constraints introduced by the exoskeleton’s structure and kinematics. Nonetheless, aside from the four isolated drops, participants were still able to complete the manipulation task with the exoskeleton.
@@
NASA-TLX scores were compared between free hand and exoskeleton conditions using one-sided Wilcoxon signed-rank tests, under the alternative hypothesis that workload would be higher with the exoskeleton (Fig. 12 and Table VI). After Bonferroni-Holm correction, the exoskeleton condition was rated significantly higher in mental demand, physical demand, effort, frustration and overall workload (all $p<.05$, one-sided). Temporal demand and performance showed no significant differences (p = .89 and .5, respectively), despite objectively measured metrics indicating otherwise.
@@
TABLE VI: NASA-TLX scores for the Six-Hole Peg Test (median [IQR]) and Wilcoxon test results (n=20).
@@
The ergonomics questionnaire results (Table IV) can be grouped into several categories. Fit (Q1–5) was rated positively, with medians between 76.5 and 85, indicating generally good adaptation of the glove and thimbles. Weight (Q6) was judged moderately heavy (Mdn = 61, IQR = 48.25–67), with most values between neutral and heavy. Restriction (Q7–9) was low overall: the index finger showed the least limitation (Mdn = 10.5), followed by the thumb (Mdn = 23), while overall hand restriction remained below 50 for almost all participants. Stability (Q10–11) was consistently high, with both fingertip and strap thimbles rated mostly stable (Mdn $\geq$ 75). Comfort (Q12, Q16) was also positive (Mdn = 67–70), although not maximal. Heat and sweating (Q13) were reported as minimal (Mdn = 2, IQR = 0–28.25). Finally, participants acknowledged an impact on perceived performance in the Six-Hole Peg test, rating completion as faster with the free hand (Q14–15, Mdn = 26–38), consistent with the observed execution times.
Overall, this first test evaluated ergonomics, mobility and the effects of wearing the exoskeleton. The device enabled most movements with minimal restriction in a group of subjects that, for the comparable parameters, represented at least approximately half of the ANSUR male population based on the most restrictive anthropometric measure, while offering satisfactory yet improvable ergonomics. With regard to movement and task performance, the Six-Hole Peg task showed increased execution time when using the exoskeleton, but very low impact on effectiveness (low failure rate).
IV-B Pick-and-place manipulation task
@@ IV-B1 Experiment definition
The second experiment assessed the influence of force feedback from the exoskeleton’s ratchet–pawl mechanism compared to free hand control in a virtual pick-and-place task. Participants were required to grasp a virtual sphere and place it into one of two target rings (Fig. 13(a)). The task was implemented in CHAI3D, which issued binary motion-blocking commands to the glove upon virtual contact detection (on: feedback engaged, off: free motion). The independent variables were:
@@
  • Tolerance threshold for failure: two values (0.025 and 0.03). During object pickup, fingertip distance was recorded and the threshold defined the allowable variation. Smaller thresholds increased the likelihood of dropping or over-squeezing the object (thus higher difficulty). These values were set after pilot testing and are not meant to represent real-world dimensions.
  • Target ring: far (Goal 1) versus near (Goal 2) from the subject, both aligned along the forward (Y) axis and equidistant from the initial ball position on the right.
@@
The hypothesis was that force feedback from the exoskeleton would reduce execution time and failure rates and yield more favorable subjective ratings, compared to free hand interaction in virtual manipulation. Objective metrics included execution time per trial, defined as the time from initial object contact to successful placement, and failures, defined as dropping or exceeding the tolerance threshold. Subjective metrics consisted of NASA-TLX for each hand condition and a questionnaire on force feedback, acoustic perception, and overall preference (Table VII).
Figure 13: Virtual setups for (a) the pick an place test and (b) perceived softness discrimination experiment.
TABLE VII: Questionnaire (Study 2)